# Beyond the Singular: The Essential Role of Multiple Generations in Effective Benchmark Evaluation and Analysis

### v2

@@ Abstract
Large language models (LLMs) have demonstrated significant utilities in real-world applications, exhibiting impressive capabilities in natural language processing and understanding. Benchmark evaluations are crucial for assessing the capabilities of LLMs as they can provide a comprehensive assessment of their strengths and weaknesses. However, current evaluation methods often overlook the inherent randomness of LLMs by employing deterministic generation strategies or relying on a single random sample, resulting in unaccounted sampling variance and unreliable benchmark score estimates. In this paper, we propose a hierarchical statistical model that provides a more comprehensive representation of the benchmarking process by incorporating both benchmark characteristics and LLM randomness. We show that leveraging multiple generations improves the accuracy of estimating the benchmark score and reduces variance. We also introduce $\mathbb{P}\left(\text{correct}\right)$, a prompt-level difficulty score based on correct ratios, providing fine-grained insights into individual prompts. Additionally, we create a data map that visualizes difficulty and semantic prompts, enabling error detection and quality control in benchmark construction.
[…]
Beyond the Singular: The Essential Role of Multiple Generations in Effective Benchmark Evaluation and Analysis
[…]
@@ 1 Introduction
There are two primary ways for generating responses from large language models (LLMs): greedy decoding and random sampling (Holtzman et al., 2019). Greedy decoding selects the next token with the highest probability, resulting in a deterministic output. In contrast, random sampling, such as nucleus sampling (Holtzman et al., 2019), incorporates randomness during decoding by sampling a token at each step based on a probability distribution. This approach leads to non-deterministic output. Current LLM benchmarks typically employ one of these methods; for instance, LiveBench (White et al., 2024) WildBench (Lin et al., 2024) and OpenLLM leaderboard (Beeching et al., 2023) use greedy decoding, while TrustLLM (Huang et al., 2024), MT Bench (Zheng et al., 2023) and Alpaca Eval (Li et al., 2023) employ a non-deterministic sampling configuration. During evaluations, LLMs generate a single response for each prompt in the benchmark, and the correctness of these responses is determined by comparing them to the ground truth answers. The final benchmark score is then calculated as the average of these individual scores.
@@ 2 Benchmarking Procedure is a Hierarchical Model
where prompt difficulty $p_{i}$ is sampled from $\mathbb{P}(\mu,\sigma;\theta)$ and $p_{i}$ represents the probability that the LLM can correctly answer the $i$-th prompt., i.e., (A generated answer to i-th prompt is $\mathbb{P}\left(\text{A generated answer to $i$-th prompt is correct}\right)=p_{i}$. This represents a latent difficulty of prompts, We denote the he $k$-th generation of the $i-$th prompt as $z_{i,j}$ and then $y_{i,j}$ is the correctness indicator for it, where $y_{i,j}=1$ if it’s correct otherwise $y_{i,k}=0$.
@@
To estimate $p_{i}$ and $\mu$, we can use a straight forward method of moment estimators $\hat{p}_{i}=\frac{\sum_{j=1}^{k}y_{i,j}}{k}$, $\hat{\mu}=\frac{\sum_{i=1}^{n}\hat{p}_{i}}{n}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{k}y_{i,j}}{nk}$. We observe that a widely used item response theory (Polo et al., 2024; Madaan et al., 2024; Ding et al., 2024), employed to model the difficulty of prompts, represents a specific parametrization of $\mathbb{P}(\mu,\sigma;\mathcal{D})$. Further elaboration on this can be found in Appendix B.
@@
Given the hierarchical model in (1) and the moment estimators
$\hat{\mu}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{k}y_{i,j}}{nk}.$
Then $\hat{\mu}$ is an unbiased estimator for $\mu$ and its variance equals:
@@
Here, $\text{Var}(\hat{\mu})$ can be decomposed into within-prompt variance and between-prompt variance. Both terms decrease as the number of benchmark data $n$ increases. However, since benchmark data is typically fixed, we analyze the influence of sampling in terms of $k$. Within-prompt variance captures the randomness in sampling $y_{ij}$ conditional on the $i-$th prompt, and it can be effectively reduced by increasing the number of samples $k$, converging to 00 as $k\to\infty$. The between-prompt variance term, on the other hand, captures the variability of prompt difficulty $p_{i}$ across groups, reflecting the randomness of difficulty distribution $\mathbb{P}(\mu,\sigma;\theta)$, and thus remains unaffected by $k$.
We can further plug in sample variance $\hat{\sigma}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(\hat{p}_{i}-\frac{\sum_{i=1}^{n}\hat{p}_{i}}{n})^{2}$ and $\hat{\mu}$ into (2) to get $\widehat{\text{Var}(\hat{\mu})}$. Finally, based on the central limit theorem, a $95\%$ confidence interval can be constructed as:
$\hat{\mu}\pm 1.96\sqrt{\widehat{\text{Var}(\hat{\mu})}}$ (3)
@@
Figure 1: Distribution of $\mathbb{P}\left(\text{correct}\right)$ of $4$ benchmarks.
@@ 2.1 Prompt Level Difficulty: $\mathbb{P}\left(\text{correct}\right)$
We refer to $\mathbb{P}\left(\text{correct}\right)=p_{i}$ in (1) and its estimation ^$\widehat{\mathbb{P}}\left(\text{correct}\right)=\hat{p}_{i}=\frac{\sum_{j=1}^{k}y_{i,j}}{k}$. When the number of generations $k$ increases, it will converge to the true $\mathbb{P}\left(\text{correct}\right)$ and therefore more fine-grained. The probability of correctness $p_{i}$ can be interpreted as a difficulty score at the prompt level: the higher the $p_{i}$, the easier the prompt since the language model has a higher probability of generating a correct response. We demonstrate the use of difficulty scores in the analysis section.
@@ 3.1 Experimental Setup
Benchmark. We choose multiple benchmarks which cover various capabilities of LLMs: MMLU-Pro (Wang et al., 2024), GSM8K (Cobbe et al., 2021), MuSR (Sprague et al., 2023), IFEval (Zhou et al., 2023). For MMLU-Pro, GSM8K, and MUSR, we use accuracy as the metric, while for IFEval, we utilize instance-level strict accuracy. Brief introduction of those benchmarks can be found in Appendix C.
@@
LLM and Setup. We utilize four widely-used open-source LLMs: Llama 3.1 (8B and 70B Instruct) (Dubey et al., 2024), Qwen 2.5 (7B Instruct) (Yang et al., 2024), and Ministral (8B Instruct) (Jiang et al., 2023)***Ministral models and analysis on Ministral output were run only by Wenbo Zhang on academic research systems.. We evaluate both greedy decoding and random sampling on these models, with the latter using a temperature of $0.7$ and top-p of $1.0$. For each prompt across all benchmarks, we generate $50$ samples ($k=50$) using a 0-shot chain-of-thought prompting strategy.
@@
Table 1: Results on four benchmark datasets with four open source LLMs. "n" is the number of prompts, "Greedy" denotes greedy decoding, "Sample (k=50)" is the random sample with $50$ generations and "$\Delta$ ($k=1$)" denotes the performance gap between the best and worst run with $1$ generation. We include both benchmark score and SE.
@@
3.2 Main Results
Our results are presented in Figures 1 and 2, as well as in Table 1. The key takeaways from these results are summarized below.
Distribution of $\mathbb{P}\left(\text{correct}\right)$ show diffuse density in challenging tasks, behaving like random samplers. For the distribution of $\mathbb{P}\left(\text{correct}\right)$, we define stable behavior as a density distribution with high concentrations near 00 and $1$, and lower density in between. Conversely, a distribution with a high density between 00 and $1$ indicates high randomness. As shown in Figure 1, when confronted with benchmarks that require strong reasoning skills(MMLU-Pro, IFEval, and MuSR), all models display a diffuse density distribution over the support $\left[0,1\right]$. This suggests that LLMs resemble random samplers when handling prompts requiring strong reasoning, underscoring the complexity and sensitivity of their reasoning processes. In contrast, the simpler task GSM8K display densities with more pronounced tails and reduced uncertainty. A plausible explanation is that GSM8K is easier and involves shorter reasoning lengths, which in turn decreases the likelihood of diverse reasoning paths emerging. Additionally, we observe that the Llama 70B model exhibits the most stable performance across all benchmarks, suggesting that larger models may be capable of providing more stable reasoning.
Estimation differs noticeably between greedy decoding and random sampling, with a single random generation being unstable. Table 1 presents the benchmark scores, highlighting the performance differences between greedy decoding and random sampling. Notably, for GSM8K and MuSR, the absolute differences in benchmark score between these two methods for Llama3 8B are $3.4$ and $4.2$ respectively, indicating a relatively large performance gap. This discrepancy can also be observed in other models and datasets. Furthermore, we observe considerable variability with one generation, characterized by large values of $\Delta$($k=1$). This suggests that random sampling with limited generations is ineffective for benchmark evaluation, particularly for small datasets, aligning with our Lemma 2. We also compared the confidence intervals with the bootstrap confidence intervals commonly used in previous studies and found similar performance. This indicates that running the bootstrap procedure is unnecessary, as our method achieves comparable results and operates much faster without resampling. We also investigate how sampling parameters influence the $\mathbb{P}\left(\text{correct}\right)$ distribution and results are in Appendix D.
Figure 2: Data map for GSM8K with Llama 70b.
Multiple generations can help detect labeling errors: a case study on GSM8K. Benchmark construction can involve label errors or ambiguous prompts, such as the approximately $5\%$ error rate in GSM8K. Manually cleaning large datasets is costly, but we found using multiple generations from advanced LLMs can help identify mislabeled or ambiguous prompts. Based on multiple generations, we can create a data map to visualize $\mathbb{P}(\text{correct})$ against $\mathbb{S}(\text{consistency})$, which measures the semantic consistency of generations. More details on this measure are provided in Appendix E. We hypothesize that prompts with low $\mathbb{P}(\text{correct})$ and high $\mathbb{S}(\text{consistency})$ may be mislabeled or ambiguous due to contradicting with the self-consistency(Wang et al., 2022). Self-consistency (Wang et al., 2022; Mitchell et al., 2022) leverages the intuition that a challenging reasoning problem typically admits multiple reasoning paths leading to its unique correct answer. To verify our hypothesis, we utilize the data map of Llama3 70B for GSM8K and selected prompts with $\mathbb{P}(\text{correct})\leq 0.1$ and $\mathbb{S}(\text{consistency})\geq-0.8$, totaling $18$ prompts. After manually reviewing the selected prompts, we found that $44.4\%$ prompts were either mislabeled or ambiguous. Examples are shown in the Appendix Figure 4. Our results demonstrate the potential of using data maps as a tool to clean datasets, which is similar to a prior study (Swayamdipta et al., 2020) but it focuses on classification models rather than generative models. Notably, our findings utilize only a single LLM and a simple semantic metric. This highlights future research that will incorporate more models and improved semantic metrics for more accurate detection.
4 Conclusion
In this paper, we investigate the role of multiple generations in enhancing LLM benchmark evaluation. Using a hierarchical model, we quantify prompt difficulty and visualize its distribution, revealing insights into benchmark data. Our findings about variance reduction, and labeling error detection emphasize the importance of leveraging multiple generations for robust benchmark evaluations.
Limitations
Ethic Statement
Acknowledgements
We thank Christine Agarwal, Xu Chen, Aude Hofleitner, Jenny Hong, Kartik Khandelwal, Wesley Lee, Tony Lee, and Yilin Zhang for their helpful discussions.
Appendix A Related Work
A.1 LLM Benchmark Evaluation
Recent benchmark evaluations have significantly enhanced our understanding of Large Language Models (LLMs) and have driven further advancements in the field. Notable benchmarks like MMLU (Hendrycks et al., 2020), HELM (Liang et al., 2022), and BIG-bench (Srivastava et al., 2022) have expanded assessments to include language generation, general knowledge understanding, and complex reasoning. Several other benchmarks assess the trustworthiness of large language models (LLMs) (Wang et al., 2023; Huang et al., 2024; Zhang et al., 2024) in terms of safety, bias, privacy, and hallucination, etc. Leaderboards like the OpenLLM Leaderboard (Beeching et al., 2023) facilitate performance comparisons across LLMs by evaluating a range of tasks, each targeting different capabilities, to provide a comprehensive assessment of LLMs. However, most benchmark evaluations, even on leaderboards, rely on a single output per example, either greedy decoding or random sampling. Song et al. (2024) also examines the performance gap between the two types of generation strategies and highlights the importance of randomness. There is also concurrent work by Miller (2024) that mentions using multiple generations to reduce variance, but their contribution is primarily conceptual. In contrast, we provide both theoretical support and empirical results. Additionally, we propose several benefits of using multiple generations, such as difficulty quantification and mislabeled prompt detection, which distinguish our work from theirs.
A.2 Prompt Difficulty in Benchmark
Understanding prompt-level difficulty is crucial for analyzing benchmark composition and some benchmark datasets include difficulty scores for each prompt provided by humans. For example, the MATH dataset (Hendrycks et al., ) offers a variety of high-school-level problems with a broad five-level difficulty rating. Similarly, the GPQA dataset (Rein et al., 2023) contains graduate-level multiple-choice questions rated on a 4-point scale by two experts. Recent studies (Ding et al., 2024; Polo et al., ) also attempted to estimate difficulty scores of individual prompts using item response theory (Cai et al., 2016; Natesan et al., 2016) or Glicko-2 (Glickman, 2012), based on offline evaluation results from a pool of large language models (LLMs) or human participants. This approach seeks to provide an objective difficulty score by encompassing a diverse range of testers, including both humans and LLMs. However, this can lead to misalignment when focusing solely on a target LLM. A question that is easy for one model might be difficult for others, highlighting the inherently subjective nature of difficulty (Desender et al., 2017). Therefore, it is more relevant to consider the subjective difficulty specific to the target LLM.
$\mathbb{P}\left(\text{correct}\right)$ is closely connected to item response theory. Many studies (Polo et al., 2024; Madaan et al., 2024; Ding et al., 2024) utilize IRT to quantify the difficulty of prompts using multiple LLMs. One variation of the IRT model is the one-parameter logistic (1PL) model as defined below:

### v3

# Beyond the Singular: Revealing the Value of Multiple Generations in Benchmark Evaluation

@@ Abstract
Large language models (LLMs) have demonstrated significant utility in real-world applications, exhibiting impressive capabilities in natural language processing and understanding. Benchmark evaluations are crucial for assessing the capabilities of LLMs as they can provide a comprehensive assessment of their strengths and weaknesses. However, current evaluation methods often overlook the inherent randomness of LLMs by employing deterministic generation strategies or relying on a single random sample, resulting in unaccounted sampling variance and unreliable benchmark score estimates. In this paper, we propose a hierarchical statistical model that provides a more comprehensive representation of the benchmarking process by incorporating both benchmark characteristics and LLM randomness. We show that leveraging multiple generations improves the accuracy of estimating the benchmark score and reduces variance. Multiple generations also allow us to define $\mathbb{P}\left(\text{correct}\right)$, a prompt-level difficulty score based on correct ratios, providing fine-grained insights into individual prompts. Additionally, we create a data map that visualizes difficulty and semantics of prompts, enabling error detection and quality control in benchmark construction.
Beyond the Singular: Revealing the Value of Multiple Generations in Benchmark Evaluation
footnotetext: * Work done during internship at Metafootnotetext: $\dagger$} Co-correspondence
@@ 1 Introduction
There are two primary ways for generating responses from large language models (LLMs): greedy decoding and random sampling (holtzman2019curious). Greedy decoding selects the next token with the highest probability, resulting in a deterministic output. In contrast, random sampling, such as nucleus sampling (holtzman2019curious), incorporates randomness during decoding by sampling a token at each step based on a probability distribution. This approach leads to non-deterministic output. Current LLM benchmarks typically employ one of these methods; for instance, LiveBench (white2024livebench) WildBench (lin2024wildbench) and OpenLLM leaderboard (open-llm-leaderboard-v1) use greedy decoding, while TrustLLM (huang2024trustllm), MT Bench (zheng2023judging) and Alpaca Eval (alpaca_eval) employ a non-deterministic sampling configuration. During evaluations, LLMs generate a single response for each prompt in the benchmark, and the correctness of these responses is determined by comparing them to the ground truth answers. The final benchmark score is then calculated as the average of these individual scores.
@@ 2 Benchmarking Procedure is a Hierarchical Model
$\begin{gathered}p_{i}\sim\mathbb{P}(\mu,\sigma;\theta)\quad\text{for }i=1,\cdots,n,\\ y_{i,j}\sim\text{Bernoulli}(p_{i})\quad\text{for }j=1,\cdots,k,\end{gathered}$ (1)
where prompt difficulty $p_{i}$ is sampled from $\mathbb{P}(\mu,\sigma;\theta)$ and $p_{i}$ represents the probability that the LLM can correctly answer the $i$-th prompt., i.e., $\mathbb{P}\left(\text{A generated answer to $i$-th prompt is correct}\right)=p_{i}$. This represents a latent difficulty of prompts, We denote the he $k$-th generation of the $i$-th prompt as $z_{i,j}$ and then $y_{i,j}$ is the correctness indicator for it, where $y_{i,j}=1$ if it’s correct otherwise $y_{i,k}=0$.
@@
To estimate $p_{i}$ and $\mu$, we can use a straight forward method of moment estimators $\hat{p}_{i}=\frac{\sum_{j=1}^{k}y_{i,j}}{k}$, $\hat{\mu}=\frac{\sum_{i=1}^{n}\hat{p}_{i}}{n}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{k}y_{i,j}}{nk}$. We observe that a widely used item response theory (polo2024tinybenchmarks; madaan2024quantifying; ding2024easy2hard), employed to model the difficulty of prompts, represents a specific parametrization of $\mathbb{P}(\mu,\sigma;\mathcal{D})$. Further elaboration on this can be found in Appendix B.
@@
Given the hierarchical model in (1) and the moment estimators $\hat{\mu}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{k}y_{i,j}}{nk}.$ Then $\hat{\mu}$ is an unbiased estimator for $\mu$ and its variance equals:
@@
[…]
We can further plug in sample variance $\hat{\sigma}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(\hat{p}_{i}-\frac{\sum_{i=1}^{n}\hat{p}_{i}}{n})^{2}$ and $\hat{\mu}$ into (2) to get $\widehat{\text{Var}(\hat{\mu})}$. Finally, based on the central limit theorem, a $95\%$ confidence interval is: $\hat{\mu}\pm 1.96\sqrt{\widehat{\text{Var}(\hat{\mu})}}$.
@@
Figure 1: Distribution of $\mathbb{P}\left(\text{correct}\right)$ of 44 benchmarks.
@@ 2.1 Prompt Level Difficulty: $\mathbb{P}\left(\text{correct}\right)$
We refer to $\mathbb{P}\left(\text{correct}\right)=p_{i}$ in (1) and its estimation $\widehat{\mathbb{P}}\left(\text{correct}\right)=\hat{p}_{i}=\frac{\sum_{j=1}^{k}y_{i,j}}{k}$. When the number of generations $k$ increases, it will converge to the true $\mathbb{P}\left(\text{correct}\right)$ and therefore more fine-grained. The probability of correctness $p_{i}$ can be interpreted as a difficulty score at the prompt level: the higher the $p_{i}$, the easier the prompt since the language model has a higher probability of generating a correct response. We demonstrate the use of difficulty scores in the analysis section.
@@ 3.1 Experimental Setup
Benchmark. We choose multiple benchmarks which cover various capabilities of LLMs: MMLU-Pro (wang2024mmlu), GSM8K (cobbe2021training), MuSR (sprague2023musr), IFEval (zhou2023instruction). For MMLU-Pro, GSM8K, and MUSR, we use accuracy as the metric, while for IFEval, we utilize instance-level strict accuracy. More details of benchmarks are in Appendix C.
@@
LLM and Setup. We utilize four widely-used open-source LLMs: Llama 3.1 (8B and 70B Instruct) (dubey2024llama), Qwen 2.5 (7B Instruct) (qwen2), and Ministral (8B Instruct) (jiang2023mistral) [^1] Ministral models and analysis on Ministral output were run only by some authors on academic research systems.. We evaluate both greedy decoding and random sampling on these models, with the latter using a temperature of $0.7$ and top-p of $1.0$. For each prompt across all benchmarks, we generate $50$ samples ($k=50$) using a 0-shot chain-of-thought prompting strategy.
@@
Table 1: Results on four benchmark datasets with four open source LLMs. "n" is the number of prompts, "Greedy" denotes greedy decoding, "Sample (k=50)" is the random sample with $50$ generations and $\Delta$ (k=1k=1)" denotes the performance gap between the best and worst run with 11 generation. We include both benchmark score and SE.
Benchmark n Llama 3.1 8b Instruct Llama3.1 70b Instruct
Greedy Sample ($k=50$) $\Delta(k=1)$ Greedy Sample ($k=50$) $\Delta(k=1)$
GSM8K $1,319$ $86.1$ ($0.95$) $85.6$ ($0.68$) $18.6$ $95.6$ ($0.56$) $95.3$ ($0.45$) $4.8$
IFEval $541$ $74.5$ ($1.87$) $71.1$ ($1.51$) $8.3$ $82.6$ ($1.64$) $80.2$ ($1.42$) $5.9$
MuSR $756$ $24.8$ ($1.65$) $29.0$ ($1.00$) $8.2$ $56.3$ ($1.80$) $57.9$ ($1.40$) $5.4$
3.2 Main Results
Results are shown in Figures 1 and Table 1, with the full table available in Appendix Table 2. Key takeaways are summarized below.
Distribution of $\mathbb{P}\left(\text{correct}\right)$ show diffuse density in challenging tasks, behaving like random samplers. For the distribution of $\mathbb{P}\left(\text{correct}\right)$, we define stable behavior as a density distribution with high concentrations near 0 and $1$, and lower density in between. Conversely, a distribution with a high density between 0 and $1$ indicates high randomness. As shown in Figure 1, when confronted with benchmarks that require strong reasoning skills (MMLU-Pro, IFEval, and MuSR), all models display a diffuse density distribution over the support $\left[0,1\right]$. This suggests that LLMs resemble random samplers when handling prompts requiring strong reasoning, underscoring the complexity and sensitivity of their reasoning processes. In contrast, the simpler task GSM8K display densities with more pronounced tails and reduced uncertainty. A plausible explanation is that GSM8K is easier and involves shorter reasoning lengths, which in turn decreases the likelihood of diverse reasoning paths emerging. Additionally, we observe that the Llama 70B model exhibits the most stable performance across all benchmarks, suggesting that larger models can provide more stable reasoning process.
Estimation differs noticeably between greedy decoding and random sampling, with a single random generation being unstable. Table 1 presents the benchmark scores, highlighting the performance differences between greedy decoding and random sampling. Notably, for GSM8K and MuSR, the absolute differences in benchmark score between these two methods for Llama3 8B are $3.4$ and $4.2$, respectively, indicating a relatively large performance gap. This discrepancy can also be observed in other models and datasets. Furthermore, we observe considerable variability with one generation, characterized by large values of $\Delta$($k=1$). This suggests that random sampling with limited generations is ineffective for benchmark evaluation, particularly for small datasets, aligning with our Lemma 2. We also investigate how sampling parameters influence the $\mathbb{P}\left(\text{correct}\right)$ distribution, and results are in Appendix D. We further conduct a synthetic analysis to demonstrate the value of multiple generations. Using $k=50$ as the oracle (i.e., the full set of generated samples), we evaluate $k=1,5,10,20$ over 1000 trials each by sampling with replacement. As shown in Fig. 2, increasing $k$ leads to narrower $95\%$ confidence intervals that coverage the true score. In contrast, greedy decoding exhibits a consistent performance gap, suggesting that even a modest number of sampled generations better approximates $\mathbb{P}(\text{correct})$ than greedy decoding.
Figure 2: Benchmark score of IFEval over different $k$.
Figure 3: Data map for GSM8K with Llama 70b.
[…]
4 Conclusion
In this paper, we investigate the value of multiple generations in LLM benchmark evaluation. By leveraging a hierarchical model, we show that multiple generations help quantify prompt difficulty, reduce variance, and detect labeling errors, making evaluations more robust and informative.
Limitations
Ethic Statement
Appendix A Related Work
A.1 LLM Benchmark Evaluation
Recent benchmark evaluations have significantly enhanced our understanding of Large Language Models (LLMs) and have driven further advancements in the field. Notable benchmarks like MMLU (hendrycks2020measuring), HELM (liang2022holistic), and BIG-bench (srivastava2022beyond) have expanded assessments to include language generation, general knowledge understanding, and complex reasoning. Several other benchmarks assess the trustworthiness of large language models (LLMs) (wang2023decodingtrust; huang2024trustllm; zhang2024defining) in terms of safety, bias, privacy, and hallucination, etc. Leaderboards like the OpenLLM Leaderboard (open-llm-leaderboard-v1) facilitate performance comparisons across LLMs by evaluating a range of tasks, each targeting different capabilities, to provide a comprehensive assessment of LLMs. However, most benchmark evaluations, even on leaderboards, rely on a single output per example, either greedy decoding or random sampling. song2024good also examines the performance gap between the two types of generation strategies and highlights the importance of randomness. There is also concurrent work by miller2024adding that mentions using multiple generations to reduce variance, but their contribution is primarily conceptual. In contrast, we provide both theoretical support and empirical results. Additionally, we propose several benefits of using multiple generations, such as difficulty quantification and mislabeled prompt detection, which distinguish our work from theirs.
A.2 Prompt Difficulty in Benchmark
Understanding prompt-level difficulty is crucial for analyzing benchmark composition and some benchmark datasets include difficulty scores for each prompt provided by humans. For example, the MATH dataset (hendrycks2measuring) offers a variety of high-school-level problems with a broad five-level difficulty rating. Similarly, the GPQA dataset (rein2023gpqa) contains graduate-level multiple-choice questions rated on a 4-point scale by two experts. Recent studies (ding2024easy2hard; polotinybenchmarks) also attempted to estimate difficulty scores of individual prompts using item response theory (cai2016item; natesan2016bayesian) or Glicko-2 (glickman2012example), based on offline evaluation results from a pool of large language models (LLMs) or human participants. This approach seeks to provide an objective difficulty score by encompassing a diverse range of testers, including both humans and LLMs. However, this can lead to misalignment when focusing solely on a target LLM. A question that is easy for one model might be difficult for others, highlighting the inherently subjective nature of difficulty (desender2017subjective). Therefore, it is more relevant to consider the subjective difficulty specific to the target LLM.
$\mathbb{P}\left(\text{correct}\right)$ is closely connected to item response theory. Many studies (polo2024tinybenchmarks; madaan2024quantifying; ding2024easy2hard) utilize IRT to quantify the difficulty of prompts using multiple LLMs. One variation of the IRT model is the one-parameter logistic (1PL) model as defined below:
Table 2: Results on four benchmark datasets with four open source LLMs. "n" is the number of prompts, "Greedy" denotes greedy decoding, "Sample (k=50)" is the random sample with $50$ generations and $\Delta$ (k=1k=1)" denotes the performance gap between the best and worst run with 11 generation. We include both benchmark score and SE.